# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 4.3 Significance level of the period
If the significance level is very close to zero, the peak could not be considered as generated by the noise and thus a high suspicion is permitted that the outstanding peak is due to a deterministic periodic signal. In the case of even sampling of the time-series, such a significance could be easily computed. However, in the case of irregular sampling and particularly in the case of the Gaia sampling, the situation is much more complex. Unfortunately, no rigorous analytical formula actually exists for these precise circumstances. Consequently, to test for the significance of the period, the procedure relies on the use of Monte-Carlo simulations performed in the conditions of the Gaia RVS sampling. In principle, this should be performed individually for each object characterised by their own sampling, but this is not affordable for such a large number of sources.
The first dependency of the statistical distribution is on the number of data points in the time-series. The secondary dependency concerns the actual distribution in time of the data points. The dispersion of the distribution decreases with increasing. Therefore, we first attempted to perform such simulations for each of the possibilities on this number . Keeping this number constant, it is easy to notice through the simulations that the statistical distributions of the highest peak at fixed remain quite similar whatever is the actual distribution in time of the data points (i.e. over the various simulations). This is particularly true for the interesting regime where the significance level (SL) is lower than 0.2, but also when is getting large (see Sect. 4.4). Therefore, various Monte-Carlo simulations were performed and the value of the height of the highest peak associated to various fixed significance levels has been computed. From these simulations, a look-up table is built that is used by the algorithm.
@@ 4.4 Build-up of the look-up table
For each source, we computed 10000 HMM periodograms of Gaussian white-noise realisations on the very same date sampling as that of the considered source, by adopting the same frequency step and limits as in the operation pipeline. For each periodogram, we retained the largest value of the , the so-called statistic. Then we merged all the values and ranked them by increasing order and build up an Empirical Cumulative Distribution Function (ECDF) with the 10000 values of the statistic. At fixed, each realisation of the sampling provides a particular ECDF function. The chosen length 10000 of the ECDF allows us to assign false alarm probabilities with a precision better than 0.5 % at 95 % of confidence.
@@
When calculating the periodogram of an input time-series with the length , its corresponding adopted ECDF (on the basis of the value) is read from the look-up table. The table has a comprehensive coverage of the number of transits, without gaps from = 10 to 72, but there are some entries for large values that are not present. In this case, the algorithm opts for the closest lowest input.
5 The calculation of the orbital solutions
@@
The Keplerian fitting, even at fixed period, is time-consuming if it is to be performed over a vast set of periods: as for example in a genuine Keplerian periodogram (see e.g. Cumming 2004; Zechmeister & Kürster 2009, their sections 2.3 and 5, respectively).
Therefore, we decided to strew subtleties over our procedure in order to render it much more robust. The first step is to search for periodicities using a periodogram (HMM method) as described in Sect. 4. This corresponds to a fit of sine/cosine, which is a linear process (at fixed frequency). This is particularly convenient: even if the RV curve is highly eccentric, and thus the power of the signal is spread over several harmonics, the power related to the period of the orbit is still dominating the noiseless periodogram. If the number of cycles is sufficient, the periodogram will always present increased power at the correct period. However, the detection of a periodicity does not imply that the motion is Keplerian. Therefore, after the establishment of a list of candidate periods, it remains to fit the Keplerian model for each of them and to deduce the orbital solution and corresponding parameters that better fit the data. We first perform a Keplerian fit that is linearised at the maximum (but not entirely) and represents a preliminary fit, not definitive but robust. The actual least-square fit is performed with the true anomaly as the independent variable.
The Fourier periodograms have the reputation to be inefficient at addressing the case of eccentric RV curves (for , see e.g. Endl et al. 2002; Cumming 2004). We tend to think that this is due to the fact that, for eccentric orbits by opposition to the circular case, not all the data points contain the same amount of information. In fact, for very eccentric orbits, the data points related to observations around the maximum separation are more determining, in particular for the semi-amplitude. If such data points are missing, have not been acquired, this will have major impact at least on the determination of the value. Consequently, this will have a huge importance on the detectability of the signal: more abruptly, we cannot detect something that is not observed. Actually, this is not a problem of ability of the Fourier periodogram to address eccentric orbits, this is a problem of fair sampling of the RV curve. The correct way to interpret things is that if a period is detected, we are authorised to have some confidence in its existence. Therefore, the nature of the Fourier periodogram improves the reliability and robustness of the algorithm at the expense of a difficulty to have a proper estimation of the completeness. This is a choice we made.
@@
6 The main NSS-SB1 processing chain
@@ 6.1 Global structure of the chain
Five Java packages operate in sequential order in our pipeline. These are InputChecker, PeriodFinder, SB1Refiner, SB1CRefiner and TrendSolution. The package InputChecker reads the input time-series in the NSS data model and copy them into arrays of primitives that are shared with the entire pipeline. This reduces the number of modifications required after every data model update. It also performs sanity checks, data cleaning and some other modifications. Finally, it passes the rearranged data to the four remaining processing packages.
@@
The second package (PeriodFinder) is working in two steps. Its first step is to establish a list of candidate periods that are significant. Its second task is to test each of the significant candidate periods by fitting the corresponding Keplerian model to the data, and to select the best period on the basis of the comparison of the various models. The best of these models represents a first approximation of the orbit. The third package (SB1Refiner) has the charge to start from the selected approximate solution and to refine it. As part of this computation, it derives the variance-covariance matrix of the fit. The fourth package (SB1CRefiner) addresses the case of circular orbits when the eccentric solution generates a badly configured variance-covariance matrix, and tries to adopt the best compromise between a circular solution and a singular eccentric one. After these computations, a decision is to be taken to adopt the most adapted periodic orbital solution. The fifth package (TrendSolution) deals with the RV time-series exhibiting a trend as a function of time (see Sect. 3.3). This means time-series that can be fitted by polynomials. These are acceleration models. After the execution of all these packages, a choice is made to select the solution that presents the best fit: SB1, SB1C, TrendSB1 or StochasticSB1.
@@
The five packages are fully described in the respective five Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5. The flow chart of the full NSS-SB1 processing is described in Fig. 4 along with the relations between the five packages mentioned above. It is interesting to notice that after the upper violet lozenge that contains a test on (that will be defined in Sect. 6.3.1), there is a bifurcation towards two sub-chains that receive the data. One sub-chain is treating the objects exhibiting a periodicity that could conduct to an orbital solution. The other sub-chain treats the cases where no periodicity is involved (only trends). From the purely logical point of view, they are working in parallel although from the practical point of view, they are executed in sequence. During the processing of the full NSS-SB1 chain, some flags are raised to retain more information about the path followed by the time-series, as well as to point out a few characteristics that could be useful for further treatment. In Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5, the flags are identified by their number and a full description of their nature is given in Sect. 6.5.
@@ 6.2 Generalities about the fits
Moreover, thanks to the eigenvectors, singular algebraic systems can still be solved by the Moore-Penrose pseudo-inverse technique (Press et al. 2007, Sect. 2.6.2). This is extremely useful because it solves the discontinuity problem between circular and eccentric models. Indeed, when eccentricity approaches zero, and parameters become progressively redundant (Sterne 1941), and thus the eccentric model becomes overstated, too powerful and thus prone to overfitting. The boundary between the two regimes, circular and eccentric, can not be set by a fixed value of the eccentricity since it is governed by the signal-to-noise ratio of the input time-series, and thus depends on many parameters such as the sampling, the amplitude of the signal and the involved noise (Lucy & Sweeney 1971).
@@ 6.3.1 Ingestion of the data
All stars that present RV variations according to the following criteria rv_chisq_pvalue0.01 and rv_normalised_gof4 are processed by the pipeline starting with the InputChecker described here. Other details about these input data were already provided in Sect. 2.
@@
NSS input data are requested from the Main Data Base (MDB) and copied into StarObject Java objects during data ingestion. A StarObject contains the single star astrometric solution (the reference epoch, the mean position, the proper motions, the parallax, ruwe and astrometric_excess_noise), the atmospheric parameters as measured by CU8 (Coordination Unit Astrophysical Parameters, Andrae et al. (2023); Recio-Blanco et al. (2023)), and the light curve classification along with the orbital period if the corresponding source is suspected to be an eclipsing binary by CU7 (Coordination Unit Variability Analysis, Mowlavi et al. (2023)). In addition to these mean data, a StarObject contains an array of Transit Java objects, where each Transit contains the relevant epoch data such as the source position, the integrated flux in the , and bands, the RV (along with its uncertainty) as well as some related quality flags. Observation dates at the centre of the astrometric field are given in BJD. With the satellite spin, an observation at the centre of the RVS CCDs is performed roughly 75 s after the observation at the centre of the astrometric field: this delay has been neglected. The observation date, the RV and its uncertainty are extracted from each Transit and copied into arrays of primitives for limiting code modifications required after every data model update. Those that are not finite numbers are filtered out. Observation dates and error bars must be positive numbers, while RVs must fall within the range of km s-1. The length of the time-series is expected to be larger than or equal to 10 and if it is not the case a flag is raised (flag nr.8, see Sect. 6.5) and the star is rejected from further processing.
@@
After these sanity checks, we start by cleaning up non-valuable entries. A Transit is discarded when its RV is detected by CU6/MTA as an outlier, either on its own or together with the rotational velocity or the integrated flux in the band or both. In addition, we also discard the Transit if it exceptionally corresponds to a composite spectrum (see Damerdji et al. 2024). The resulting time-series after cleaning contains data points distributed over a genuine individual time span (the difference between the two dates of observations at the extreme of the interval).
@@
Finally, we perform a few modifications. We correct uncertainties for sources with large ruwe, as described in Sect 2. The reference epoch J2016.0 (BJD 2457389.0) of the astrometric solutions, is subtracted from the observation dates for reducing the correlations between the derived solution parameters. A cleaned time-series of RVs, still containing the required number of 10 observations, is shared with the whole software for the rest of processing, whilst, in the opposite case, it is rejected from the pipeline, and the corresponding flag is raised (flag nr.10, see Sect. 6.5 for flag descriptions). The rejection from the pipeline is done by outputting an empty solution containing the flag and no parameter values. Such an empty solution is for internal use and statistical analysis. It is not further considered for the final output of the chain.
@@ 6.3.2 Inventory of the periods and approximate model
The package PeriodFinder is first establishing a list of candidate frequencies. The time-series of length is analysed thanks to an HMM periodogram, as explained in Sect. 4.2. The package reads the ECDF corresponding to (see Sect. 4.4) and establishes a threshold above which the peak is considered significant against the null-hypothesis of white-noise. All the peaks in the periodogram that are above the threshold value have their position in frequency entering the list. This is clearly an approximation since the test is constructed on the rigorous concept that we are looking at the highest peak only (see Sect. 4.3). No fully sound counter-indication to this assumption has been identified yet. Up to 100 frequencies/periods are accepted in the list. No additional smaller peak is considered above this amount. If the object under analysis is an Eclipsing Binary, we also add to the list the value of the period issued from the photometry. We test the significance of any available external period (i.e. in DR3 from CU7) before adding it to the set of candidate periods. For DR3 the periods entering the list were selected on the basis of an adopted threshold probability value conf_spectro_period larger than 0.95 ( smaller than 0.05). If the list of significant candidate periods is empty, no full orbital solution is possible. The relevant flag is raised (flag nr.13).
@@
where and . From these and , the Campbell parameters and are easily extracted. In Equation (10), , , and appear in a linear way. They can be optimised through an ordinary linear least-square regression and this for every set of trial . This system of three linear and three non-linear parameters, can be efficiently solved with the separable variable technique (Golub & Pereyra 2003). Indeed, we optimise three non-linear parameters only: , and (1/). Consequently, the dimensional space of variables is reduced, leading to more robust and rapid convergences. The pseudo-inverse implemented in the linear algebra toolbox smooths singularities near zero eccentricities. The rank of the involved linear systems is smaller than or equal to three, allowing them to be solved using more precise analytical computations. We give in Appendix C a summary of the involved computations. The minimisation is limited to 1000 iterations for each of the 8 considered starting points, and this for every candidate frequency . The solution is taken into account for further comparison, even when the maximum number of iterations is reached. The relative convergence of the 8 different descending tracks are analysed to estimate the robustness of the convergence (the arrival point should be the same in a perfect world). Ultimately, the best solution is selected for the considered candidate frequency. This algorithm is quite robust and provides adequate fits for orbital curves up to eccentricities of 0.995, well beyond what is necessary for the majority of the objects. The procedure is executed for every candidate frequency and gives a of the fit that is translated into an statistic. The smallest value identifies the best fit and the best set of parameters, including the period. The algorithm is efficient at finding the best minimum, but the fit is performed in two steps which is clearly an approximation concerning the cross-covariance terms between families of parameters (linear versus non-linear). Therefore, we need to refine the approximate model.

### v2

@@ 4.3 Significance level of the period
If the significance level is very close to zero, the peak could not be considered as generated by the noise; thus, there is considerable suspicion that the outstanding peak is due to a deterministic periodic signal. In the case of even sampling of the time series, such a significance could be easily computed. However, in the case of irregular sampling and particularly in the case of the Gaia sampling, the situation is much more complex. Unfortunately, no rigorous analytical formula actually exists for these precise circumstances. Consequently, to test for the significance of the period, the procedure relies on the use of Monte-Carlo simulations performed in the conditions of the Gaia RVS sampling. In principle, this should be performed individually for each object characterised by their own sampling, but this is not affordable for such a large number of sources.
The first dependency of the statistical distribution is on the number of data points in the time series. The secondary dependency concerns the actual distribution in time of the data points. The dispersion of the distribution decreases with increasing. Therefore, we first attempted to perform such simulations for each of the possibilities on this number . Keeping this number constant, it is easy to notice through the simulations that the statistical distributions of the highest peak at fixed remain quite similar whatever is the actual distribution in time of the data points (i.e. over the various simulations). This is particularly true for the interesting regime where the significance level (SL) is lower than 0.2, but also when is getting large (see Sect. 4.4). Therefore, various Monte-Carlo simulations were performed and the value of the height of the highest peak associated to various fixed significance levels has been computed. From these simulations, a look-up table is built that is used by the algorithm.
@@ 4.4 Build-up of the look-up table
For each source, we computed 10 000 HMM periodograms of Gaussian white-noise realisations on the very same date sampling as that of the considered source, by adopting the same frequency step and limits as in the operation pipeline. For each periodogram, we retained the largest value of the , the so-called statistic. Then we merged all the values and ranked them by increasing order and build up an empirical cumulative distribution function (ECDF) with the 10 000 values of the statistic. At fixed, each realisation of the sampling provides a particular ECDF function. The choice of 10 000 in the framework of the ECDF allows us to assign false alarm probabilities with a precision better than 0.5 % at 95 % confidence.
@@
When calculating the periodogram of an input time series with the length , its corresponding adopted ECDF (on the basis of the value) is read from the look-up table. The table has a comprehensive coverage of the number of transits, without gaps from = 10 to 72, but there are some entries for large values that are not present. In this case, the algorithm opts for the closest lowest input.
5 Calculating the orbital solutions
@@
The Keplerian fitting, even at fixed period, is time-consuming if it is to be performed over a vast set of periods: as for example in a genuine Keplerian periodogram (see e.g. Cumming 2004; Zechmeister & Kürster 2009, their Sects. 2.3 and 5, respectively).
Therefore, we decided to strew subtleties over our procedure in order to render it much more robust. The first step is to search for periodicities using a periodogram (HMM method) as described in Sect. 4. This corresponds to a fit of sine/cosine, which is a linear process (at fixed frequency). This is particularly convenient: even if the RV curve is highly eccentric, and thus the power of the signal is spread over several harmonics, the power related to the period of the orbit is still dominating the noiseless periodogram. If the number of cycles is sufficient, the periodogram will always present increased power at the correct period. However, the detection of a periodicity does not imply that the motion is Keplerian. Therefore, after the establishment of a list of candidate periods, the Keplerian model needs to be fit for each of them to deduce the orbital solution and corresponding parameters that are a better fit to the data. We first performed a Keplerian fit that is linearised at the maximum (but not entirely) and represents a preliminary fit, not definitive but robust. The actual least-square fit was performed with the true anomaly as the independent variable.
The Fourier periodograms are known to be inefficient at addressing the case of eccentric RV curves (for , see e.g. Endl et al. 2002; Cumming 2004). We tend to think that this is due to the fact that for eccentric orbits by opposition to the circular case, not all the data points contain the same amount of information. In fact, for very eccentric orbits, the data points related to observations around the maximum separation are more determining, in particular for the semi-amplitude. If such data points are missing, have not been acquired, this will have major impact at least on the determination of the value. Consequently, this will have a huge importance on the detectability of the signal: more abruptly, we cannot detect something that is not observed. Actually, this is not a problem of ability of the Fourier periodogram to address eccentric orbits, this is a problem of fair sampling of the RV curve. The correct way to interpret things is that if a period is detected, we are authorised to have some confidence in its existence. Therefore, the nature of the Fourier periodogram improves the reliability and robustness of the algorithm at the expense of a difficulty to have a proper estimation of the completeness. This is a choice we made.
@@
6 Main NSS-SB1 processing chain
@@ 6.1 Global structure of the chain
Five Java packages operate in sequential order in our pipeline. These are InputChecker, PeriodFinder, SB1Refiner, SB1CRefiner and TrendSolution. The package InputChecker reads the input time series in the NSS data model and copies them into arrays of primitives that are shared with the entire pipeline. This reduces the number of modifications required after every data model update. It also performs sanity checks, data cleaning and some other modifications. Finally, it passes the rearranged data to the four remaining processing packages.
@@
The second package (PeriodFinder) is working in two steps. Its first step is to establish a list of candidate periods that are significant. Its second task is to test each of the significant candidate periods by fitting the corresponding Keplerian model to the data, and to select the best period on the basis of the comparison of the various models. The best of these models represents a first approximation of the orbit. The third package (SB1Refiner) has the charge to start from the selected approximate solution and to refine it. As part of this computation, it derives the variance-covariance matrix of the fit. The fourth package (SB1CRefiner) addresses the case of circular orbits when the eccentric solution generates a badly configured variance-covariance matrix, and tries to adopt the best compromise between a circular solution and a singular eccentric one. After these computations, a decision is to be taken to adopt the most adapted periodic orbital solution. The fifth package (TrendSolution) deals with the RV time series exhibiting a trend as a function of time (see Sect. 3.3). This means time series that can be fitted by polynomials. These are acceleration models. After the execution of all these packages, a choice is made to select the solution that presents the best fit: SB1, SB1C, TrendSB1, or StochasticSB1.
@@
The five packages are fully described in the respective five Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4, and 6.3.5. The flow chart of the full NSS-SB1 processing is described in Fig. 4, along with the relations between the five packages mentioned above. It is interesting to notice that after the upper violet lozenge that contains a test on (defined in Sect. 6.3.1), there is a bifurcation towards two sub-chains that receive the data. One sub-chain is treating the objects exhibiting a periodicity that could conduct to an orbital solution. The other sub-chain treats the cases where no periodicity is involved (only trends). From the purely logical point of view, they are working in parallel although from the practical point of view, they are executed in sequence. During the processing of the full NSS-SB1 chain, some flags are raised to retain more information about the path followed by the time series, as well as to point out a few characteristics that could be useful for further treatment. In Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4, and 6.3.5, the flags are identified by their number and a full description of their nature is given in Sect. 6.5.
@@ 6.2 Generalities about the fits
Moreover, thanks to the eigenvectors, singular algebraic systems can still be solved by the Moore-Penrose pseudo-inverse technique (Press et al. 2007, Sect. 2.6.2). This is extremely useful because it solves the discontinuity problem between circular and eccentric models. Indeed, when eccentricity approaches zero, and parameters become progressively redundant (Sterne 1941), and thus the eccentric model becomes overstated, too powerful and thus prone to overfitting. The boundary between the two regimes, circular and eccentric, cannot be set by a fixed value of the eccentricity since it is governed by the signal-to-noise ratio of the input time series, and thus depends on many parameters such as the sampling, the amplitude of the signal and the involved noise (Lucy & Sweeney 1971).
@@ 6.3.1 Ingestion of the data
All stars that present RV variations according to the following criteria rv_chisq_pvalue0.01 and rv_renormalised_gof4 are processed by the pipeline starting with the InputChecker described here. Other details about these input data were already provided in Sect. 2.
@@
In the ingestion process, the NSS input data are requested from the Main Data Base (MDB) and copied into StarObject Java objects during data ingestion. A StarObject contains the single star astrometric solution (the reference epoch, the mean position, the proper motions, the parallax, ruwe, and astrometric_excess_noise), the atmospheric parameters as measured by CU8 (Coordination Unit Astrophysical Parameters, Andrae et al. (2023); Recio-Blanco et al. (2023)), as well as the light curve classification and the orbital period if the corresponding source is suspected to be an eclipsing binary by CU7 (Coordination Unit Variability Analysis, Mowlavi et al. (2023)). In addition to these mean data, a StarObject contains an array of Transit Java objects, where each Transit contains the relevant epoch data such as the source position, the integrated flux in the , and bands, the RV (along with its uncertainty) as well as some related quality flags. Observation dates at the centre of the astrometric field are given in BJD. With the satellite spin, an observation at the centre of the RVS CCDs is performed roughly 75 s after the observation at the centre of the astrometric field: this delay has been neglected. The observation date, the RV and its uncertainty are extracted from each Transit and copied into arrays of primitives for limiting code modifications required after every data model update. Those that are not finite numbers are filtered out. Observation dates and error bars must be positive numbers, while RVs must fall within the range of km s-1. The length of the time series is expected to be larger than or equal to 10 and if it is not the case a flag is raised (flag nr.8; see Sect. 6.5) and the star is rejected from further processing.
@@
After these sanity checks, we start by cleaning up non-valuable entries. A Transit is discarded when its RV is detected by CU6/MTA as an outlier, either on its own or together with the rotational velocity or the integrated flux in the band or both. In addition, we also discard the Transit if it exceptionally corresponds to a composite spectrum (see Damerdji et al. 2024). The resulting time series after cleaning contains data points distributed over a genuine individual time span (the difference between the two dates of observations at the extreme of the interval).
@@
Finally, we perform a few modifications. We correct uncertainties for sources with large ruwe, as described in Sect 2. The reference epoch J2016.0 (BJD 2 457 389.0) of the astrometric solutions, is subtracted from the observation dates for reducing the correlations between the derived solution parameters. A cleaned time series of RVs, still containing the required number of 10 observations, is shared with the whole software for the rest of processing, whilst, in the opposite case, it is rejected from the pipeline, and the corresponding flag is raised (flag nr.10; see Sect. 6.5 for flag descriptions). The rejection from the pipeline is done by outputting an empty solution containing the flag and no parameter values. Such an empty solution is for internal use and statistical analysis. It is not further considered for the final output of the chain.
@@ 6.3.2 Inventory of the periods and approximate model
The package PeriodFinder is first establishing a list of candidate frequencies. The time series of length is analysed thanks to an HMM periodogram, as explained in Sect. 4.2. The package reads the ECDF corresponding to (see Sect. 4.4) and establishes a threshold above which the peak is considered significant against the null-hypothesis of white-noise. All the peaks in the periodogram that are above the threshold value have their position in frequency entering the list. This is clearly an approximation since the test is constructed on the rigorous concept that we are looking at the highest peak only (see Sect. 4.3). No fully sound counter-indication to this assumption has been identified yet. Up to 100 frequencies/periods are accepted in the list. No additional smaller peak is considered above this amount. If the object under analysis is an Eclipsing Binary, we also add to the list the value of the period issued from the photometry. We test the significance of any available external period (i.e. in DR3 from CU7) before adding it to the set of candidate periods. For DR3 the periods entering the list were selected on the basis of an adopted threshold probability value conf_spectro_period larger than 0.95 ( smaller than 0.05). If the list of significant candidate periods is empty, no full orbital solution is possible. The relevant flag is raised (flag nr.13).
@@
where and . From these and , the Campbell parameters and are easily extracted. In Eq. (10), , , and appear in a linear way. They can be optimised through an ordinary linear least-square regression and this for every set of trial . This system of three linear and three non-linear parameters, can be efficiently solved with the separable variable technique (Golub & Pereyra 2003). Indeed, we optimise three non-linear parameters only: , and (1/). Consequently, the dimensional space of variables is reduced, leading to more robust and rapid convergences. The pseudo-inverse implemented in the linear algebra toolbox smooths singularities near zero eccentricities. The rank of the involved linear systems is smaller than or equal to three, allowing them to be solved using more precise analytical computations. We give in Appendix C a summary of the involved computations. The minimisation is limited to 1000 iterations for each of the 8 considered starting points, and this for every candidate frequency . The solution is taken into account for further comparison, even when the maximum number of iterations is reached. The relative convergence of the 8 different descending tracks are analysed to estimate the robustness of the convergence (the arrival point should be the same in a perfect world). Ultimately, the best solution is selected for the considered candidate frequency. This algorithm is quite robust and provides adequate fits for orbital curves up to eccentricities of 0.995, well beyond what is necessary for the majority of the objects. The procedure is executed for every candidate frequency and gives a of the fit that is translated into an statistic. The smallest value identifies the best fit and the best set of parameters, including the period. The algorithm is efficient at finding the best minimum, but the fit is performed in two steps, which is clearly an approximation concerning the cross-covariance terms between families of parameters (linear versus non-linear). Therefore, we need to refine the approximate model.